OpenStudy (anonymous):

Solve 3|2x + 2| - 7 = 5 for x. Show your work to receive full credit.

OpenStudy (anonymous):

3|2x + 2| = 12 |2x + 2| = 4 Then we split the equation. Part 1: 2x + 2 = 4 2x = 2 x = 1 Part 2: 2x + 2 = -4 2x = -6 x = -3 So, x = 1, -3